# Pairwise Independence of Representation, Classification, and Composition in Finite Extensional Magmas

### 27 Mar 2026 v1

@@ 10. Conclusion
Two obstructions, K-infinity and no-associativity, rule out both combinatory algebras and the classical algebraic hierarchy as settings for finite algebras combining encoding with classification. In the remaining landscape of finite extensional 2-pointed magmas, we have identified three natural algebraic properties, self-representation (R), the classifier dichotomy (D), and the Internal Composition Property (H), and proved they are pairwise independent. The six non-implications are machine-checked in Lean 4. The optimal coexistence witness has N=5N\!=\!5 (Theorem 7.2). The three-category decomposition induced by D is a functorial invariant (Theorem 4.2), and the ICP is logically equivalent to Compose+Inert (Theorem 5.1). All results are formalized in Lean 4 (80 theorems across 10 files, zero sorry).
@@
    Categorical characterization of H: Is the ICP equivalent to a standard universal property (e.g., existence of a partial internal hom)?
@@
    Categorical characterization of D: The classifier dichotomy captures the Boolean/decidability aspect of a subobject classifier (purity of classification, no intermediate truth values) but not the universality aspect (unique characteristic morphism for every subobject). The left-multiplication maps lack composition closure, precluding standard categorical formulations. Does D correspond to a known concept in non-associative categorical algebra, or does it constitute a new structural property?
  4. (4)
    Internal self-simulation: Does there exist a finite extensional 2-pointed magma satisfying R, D, and H (with the Branch and fixed-point enrichments) that supports non-trivial self-simulation: an internal evaluator that computes its own Cayley table, or a non-trivial fragment thereof, entirely from within the magma? Non-trivial self-simulation in the PCA or Turing category sense requires a fixed evaluation mechanism and universality over an open-ended domain, both absent in the finite total setting. Whether a meaningful intermediate definition exists is open.
@@
All Lean files, counterexample tables, SAT reproduction scripts, and the 16-element reflective tower artifact are available at https://github.com/stefanopalmieri/finite-magma-independence.
@@ Appendix A Cayley Tables
N=5N\!=\!5: R+D+H Optimal Coexistence ($s=r$, self-simulates).
Roles: $0=z_{1}$, $1=z_{2}$, $2=s=r$ (identity on core, ICP $b$), $3=\tau$ (classifier, ICP $a$, self-sim term), $4=\tau_{2}$ (classifier, ICP $c$).  Lean: Witness5.lean, SelfSim5.lean.
@@
N=5N\!=\!5: $D\not\Rightarrow R$ Counterexample.
No retraction pair exists. Element 2 is a classifier: row on core is $(1,0,0)\subseteq\{z_{1},z_{2}\}$. Elements 3 and 4 are non-classifiers: all core outputs $\geq 2$. Dichotomy holds; no pair of core elements forms a retraction pair.  Lean: E2PM.lean (dNoS_e2pm).
@@
2 0 0 1 0 0
3 4 2 4 2 4
4 2 2 3 3 3
N=6N\!=\!6: $H\not\Rightarrow R$ Counterexample.
No retraction pair exists. ICP holds: a=2,b=4,c=5a\!=\!2,b\!=\!4,c\!=\!5 with $b$ core-preserving and $2\cdot x=5\cdot(4\cdot x)$ on core.  Lean: E2PM.lean (hNoS_e2pm).
$\cdot$ 0 1 2 3 4 5
0 0 0 0 0 0 0
1 1 1 1 1 1 1
2 3 3 2 2 2 4
3 0 2 2 2 4 0
4 4 3 2 2 4 5
5 0 0 2 2 2 4
@@
R+D+H roles: $0=z_{1}$, $1=z_{2}$, $2=s$, $3=r$, $4=\tau$, $8=\eta$ (ICP $a$), $6=g$ (ICP $b$, core-preserving), $7=\rho$ (ICP $c$). Enrichment roles (not used in this paper’s theorems): $5=f$, $7=\rho$ (Branch dispatch), $9=Y$ (fixed point).  Lean: Witness10.lean.
@@ Appendix B Proof Inventory
Functoriality 4 Algebraic Invariance under iso
@@
E2PM 10 decide $D\not\Rightarrow R$, $H\not\Rightarrow R$, $R\not\Rightarrow H$

### 04 Apr 2026 v2

@@ 10. Conclusion
Two obstructions, K-infinity and no-associativity, rule out both combinatory algebras and the classical algebraic hierarchy as settings for finite algebras combining encoding with classification. In the remaining landscape of finite extensional 2-pointed magmas, we have identified three natural algebraic properties, self-representation (R), the classifier dichotomy (D), and the Internal Composition Property (H), and proved they are pairwise independent. The six non-implications are machine-checked in Lean 4. The optimal coexistence witness has N=5N\!=\!5 (Theorem 7.2). The three-category decomposition induced by D is a functorial invariant (Theorem 4.2), and the ICP is logically equivalent to Compose+Inert (Theorem 5.1). All results are formalized in Lean 4 (93 theorems across 12 files, zero sorry).
@@
    Categorical characterization of H: The ICP is the weakest non-degenerate factorization property of the left-regular representation (weakening either pairwise-distinctness or non-triviality makes it vacuous). Is it equivalent to a standard universal property (e.g., existence of a partial internal hom)?
@@
    Categorical characterization of D: D is the unique coarsest non-trivial isomorphism-invariant partition of core (Corollary 4.3), isolating the two-valued aspect of a subobject classifier without the universality aspect. The left-multiplication maps lack composition closure, precluding standard categorical formulations. Does D correspond to a known concept in non-associative categorical algebra, or does it constitute a new structural property?
@@
All Lean files, counterexample tables, and SAT reproduction scripts are available at https://github.com/stefanopalmieri/finite-magma-independence.
@@ Appendix A Cayley Tables
N=5N\!=\!5: R+D+H Optimal Coexistence ($s=r$, minimal coexistence witness).
Roles: $0=z_{1}$, $1=z_{2}$, $2=s=r$ (identity on core, ICP $b$), $3=\tau$ (classifier, ICP $a$), $4=\tau_{2}$ (classifier, ICP $c$).  Lean: Witness5.lean.
@@
N=4N\!=\!4: $D\not\Rightarrow R$ Counterexample (tight).
Element 2 is a classifier: row on core is $(1,1)\subseteq\{z_{1},z_{2}\}$. Element 3 is a non-classifier: row on core is $(2,2)\subseteq\mathrm{core}$. No retraction pair exists among the 4 candidate $(s,r)$ assignments. Tight: D requires $N\geq 4$.  Lean: E2PM.lean (d_not_implies_s_tight).
$\cdot$ 0 1 2 3
0 0 0 0 0
1 1 1 1 1
2 0 1 1 1
3 2 3 2 2
N=5N\!=\!5: $H\not\Rightarrow R$ Counterexample (tight).
ICP holds but no retraction pair exists. Tight: ICP requires $N\geq 5$.  Lean: E2PM.lean (h_not_implies_s_tight).
@@
2 3 1 0 3 1
3 2 4 3 4 2
4 2 2 1 0 3
N=5N\!=\!5: $H\not\Rightarrow D$ Counterexample (tight).
ICP holds but the dichotomy fails: no element maps core entirely to $\{z_{1},z_{2}\}$ (no classifier exists). Tight: ICP requires $N\geq 5$.  Lean: E2PM.lean (h_not_implies_d_tight).
$\cdot$ 0 1 2 3 4
0 0 0 0 0 0
1 1 1 1 1 1
2 3 3 4 3 3
3 2 4 4 4 3
4 2 2 2 4 4
@@
R+D+H roles: $0=z_{1}$, $1=z_{2}$, $2=s$, $3=r$, $4=\tau$, $8=\eta$ (ICP $a$), $6=g$ (ICP $b$, core-preserving), $7=\rho$ (ICP $c$).  Lean: Witness10.lean.
@@ Appendix B Proof Inventory
Functoriality 4 Algebraic Decomposition invariance
CapabilityInvariance 4 Algebraic R, D, H each invariant
@@
E2PM 19 decide $D\not\Rightarrow R$, $H\not\Rightarrow R$, $H\not\Rightarrow D$, $R\not\Rightarrow H$ (tight)